together richard and and his brother frank can paint a garage in 2 days. if richard works twice as fats as frank, how long would it take each one to paint the garage alone

Answer by ptaylor(2198)

Let x=amount of time it takes Richard working alone
Then Richard works at the rate of 1/x garage per day
And let 2x=amount of time it takes Frank working alone
Then Frank works at the rate of 1/2x garage per day
We are told that together Richard and Frank work at the rate of 1/2 garage per day, so:
1/x+1/2x =1/2 multiply each term by 2x
2+1=x
x=3 days---- amount of time it takes Richard working alone
Then 2x=2*3=6 days---amount of time it takes Frank working alone
CK
1/3 +1/6 = 1/2
2/6 + 1/6=1/2
1/2=1/2
Hope this helps---ptaylor